OpenStudy (anonymous):

I don't understand this function notation, can you please help? If f(x) = x 2 + 1 and g(x) = 3x + 1, find f(-1) + 5g(-3). -40 -38 7

OpenStudy (anonymous):

are these your functions: \(\large f(x)=x^2+1 \) and \(\large g(x)=3x+1 \) ???

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

find f(-1), which x=-1, and do the same for g(x)

OpenStudy (anonymous):

f(-1) means to evaluate the expression when x=-1.. so \(\large f(-1)=(-1)^2+1 \)

OpenStudy (anonymous):

can you do that with g((-3)...????

OpenStudy (anonymous):

okay.. Let me try that...

OpenStudy (anonymous):

So, If I'm doing this right... -1 + 15 (-3) is what I solve, right?

OpenStudy (anonymous):

is that for g(-3)???

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

\(\large g(-3)=3(-3)+1 = \) ???? can you evaluate g(-3) now?

OpenStudy (anonymous):

so g(-3) would be -8

OpenStudy (anonymous):

ok.. good...

OpenStudy (anonymous):

and what was f(-1) = ????

OpenStudy (anonymous):

Dont forget the 5 in front of 5g(-3) =0

OpenStudy (anonymous):

I think that my problem is trying to get to that final step.. I can't figure out what's wrong with my work... oh and f(-1) = 1

OpenStudy (anonymous):

i don't think that's right... try again... f(-1) = ???

OpenStudy (anonymous):

well if it's (-1) squared + 1... -1 squared is 1 + positive one... oh wait.. 0??

OpenStudy (anonymous):

\(\large f(-1)=(-1)^2+1=(-1)(-1)+1=2 \)

OpenStudy (anonymous):

f(x) = x 2 + 1 and g(x) = 3x + 1, find f(-1) + 5g(-3). f(-1)=(-1)^2+1 f(-1)=1+1 f(-1)=2 5g(-3)=5*((3*(-3))+1) 5g(-3)=5*(-8) 5g(-3)=-40 f(-1)+5g(-3)=2+(-40)=-38

OpenStudy (anonymous):

I'm following..

OpenStudy (anonymous):

looks like he ^^^ did the rest for u....